i am having trouble on ways to add integers like:
4+x=2
6+n=4

Answer by Theo(13342)

4 + x = 2
subtract 4 from both sides of the equation to get:
4 + x - 4 = 2 - 4
simplify to get:
x = -2
second problem:
6 + n = 4
subtract 6 from both sides of the equation to get:
6 + n - 6 = 4 - 6
simplify to get:
n = 4 - 6
simplify to get:
n = -2
-----
when you are solving equations, you can do the following:
add the same amount to both sides of the equation.
subtract the same amount from both sides of the equation.
multiply both sides of the equation by the same amount.
divide both sides of the equation by the same amount.
-----
doing any of these thing will not change the equality of the equation.
if the left side was equal to the right side before the operation, it will still be equal to the right side after the operation.
-----
you do these things to isolate the variable that you want to solve for on one side of the equation.
usually it's the left side of the equation but doesn't have to be.
once the variable has been completely isolated to one side of the equation it has been solved.
you were left with x = -2 which means that x was solved for.
you were left with n = -2 which means that n was solved for.